# Learning Neural Contracting Dynamics: Extended Linearization and Global Guarantees

### v2

@@ 3.3 Latent Space Representation
3.4 On the Interdependence of Diffeomorphism and Learned Dynamics
To demonstrate the need for a coordinate transform, we consider the task of learning a vector field that fits trajectories generated by the linear system $\dot{x}=Ax$ with
$A=\begin{pmatrix}-1&4\\ 0&-1\end{pmatrix}.$ (17)
Clearly, this linear system cannot be represented in the form (10) with parametrization (11). To see this fact, we observe that $A+A^{\top}$ is not negative definite and thus no choice of $P_{s}$ or $P_{a}$ can represent this linear system. To remedy this issue, one can take the linear coordinate transform $z=\phi(x)=Px$, where $P=\begin{pmatrix}1&0\\ 0&4\end{pmatrix}$. Then the $z$-dynamics read
$\dot{z}=PAP^{-1}z=\begin{pmatrix}-1&1\\ 0&-1\end{pmatrix}z$ (18)
and now the symmetric part of $PAP^{-1}$ is negative definite and thus we can find suitable choices of $P_{s},P_{a}$. Specifically, we can take $\alpha\in(0,1/2)$, let $P_{a}(z)=\begin{pmatrix}0&0\\ 1/2&0\end{pmatrix}z$, and let $P_{s}(z)=Qz$ where $Q$ is the matrix square root of $\begin{pmatrix}1&-1/2\\ -1/2&1\end{pmatrix}-\alpha I_{2}$. Note that in the toy example (17), asymmetry is essential to exactly represent these dynamics in the latent space, $z$. If we used the parametrization in (Beik-Mohammadi et al., 2024), we would not be able to represent these dynamics since they have an asymmetric Jacobian.
@@ 3.5 Choice of Diffeomorphism
There are several popular neural network diffeopmorphisms including coupling layers (Dinh et al., 2016; Rana et al., 2020), normalizing flows (Urain et al., 2020), $\mathcal{M}$-flow (Brehmer & Cranmer, 2020; Beik-Mohammadi et al., 2024), spline flows (Durkan et al., 2019b), and radial basis functions (Beik-Mohammadi et al., 2023).
A coupling layer $\phi:^{d}\to^{d}$ consists of a neural network $\theta:^{k}\to^{N}$ with $1<k<d$ and a neural network $g:^{N}\times\to$. The transform $\phi$ maps input $x\in^{d}$ to $y\in^{d}$ with the following procedure:
@@
Polynomial spline curves are another diffeomorphism that are commonly used as $g$ in coupling layers. A polynomial spline has a restricted domain which is divided into bins. A different linear, quadratic (Müller et al., 2019), or cubic (Durkan et al., 2019a) polynomial covers each bin. Each polynomial is constructed so $g$ is monotonically increasing and continuous on its domain. (Durkan et al., 2019b) introduce rational-quadratic splines, which constructs a spline from functions that are the quotient of two quadratic polynomials.
In practice, $\phi$ is the composition of several of transforms. Because the coupling transform only alters some coordinates, they are often used in conjunction with random permutations. If the latent space dimension is smaller than the data space dimension (Beik-Mohammadi et al., 2024) makes the last composite function of $\phi$ an ’Unpad’ function, which simply removes dimensions to match the latent space. The decoder $\phi$ is then prepended by a ’Pad’ function, which concatenates the latent variable with the necessary number of ’zeros’ to match the data space. Hence, the coordinate transform in (Beik-Mohammadi et al., 2024) takes the form:
$\phi=\text{Unpad}\circ g_{m}\circ\dots\circ g_{1}$ (19)
$\phi^{-1}=g_{1}^{-1}\circ\dots\circ g_{1}^{-1}\circ\text{Pad}$ (20)
where $g_{i}$ are all either coupling transforms or permutations.
Of course, if the latent dimension is smaller than that of the data dimension, then $\phi$ can not be bijective. However, (Beik-Mohammadi et al., 2024) argue that as long as $\phi$ is injective and has range over the dataset, then $f$ being contracting in the latent space implies that the learned dynamics are contracting on the submanifold defined by the image of $\phi^{-1}$. In other words, NCDS cannot be globally contracting. The consequences of this are shown in Figure (3). While NCDS can learn to admit an equilibrium point when on the submanifold, trajectories that fall off the submanifold may diverge. This is especially prevalent during early phases of training. ELCD, in contrast, is contracting to the correct equilibrium point during all phases of training.
3.6 Training
Our task is to simultaneously learn the contracting system $f(x)$ and the metric in which $f$ contracts. As previously discussed, the metric is implicitly determined by $\phi(x)$. (Beik-Mohammadi et al., 2024) uses a two-step method. First, they treat $\phi$ as a variational autoencoder and minimize the evidence lower bound (ELBO). They then fix the encoder and train the model to evolve the state in latent space. Note that the VAE objective is to make the encoded data to match a standard-normal distribution. Notably, VAE training does not encourage the encoder to transform the data to space in which the data is corresponds to trajectories of a contracting system. If the data is not contracting in the transformed space, a contracting model will not be able to fully fit the data. This explains why, in our implementation of the two step-training, we are unable to reasonably learn the data. For further comparison with NCDS, we will train the encoder and model jointly.
@@ 3.6 Training
Figure 3: Plots of the vector fields and induced trajectories learned by ELCD (Top) and NCDS (Bottom) after different training epochs. ELCD is guaranteed to contract to a given equilibrium point even when untrained, and will do so over all training epochs. NCDS, in contrast, does not parameterize its equilibrium point explicitly, and so it must be learned. In early training epochs, learned trajectories can veer outside the data as shown above.
@@ 4.1 Datasets
Evaluating contracting neural networks is difficult, as the task of learning contracting dynamical systems does not align perfectly with learning noisy demonstration trajectories. Nevertheless, we experiment with the LASA dataset (Lemme et al., 2015), which consists of 30, two-dimensional curves. We use three demonstration trajectories of each curve. As in (Beik-Mohammadi et al., 2024), the first few initial points of each trajectory are omitted so that only the target state has zero velocity. we stack two and four LASA trajectories together to make datasets of 4 and 8-dimensional trajectories, respectively. All data is standardized to have a mean of zero and variance of one. We us in total 10 2D curves, 3 4D curves, and 2 8D curves. The 2D-LASA trajectories and their respective trained ELCD trajectories and vector fields are visualized in Figure (2).
4.2 Metric
@@
We evaluate trajectories by comparing them against their reference trajectory using the normalized dynamic time warping distance (DTWD) (Keogh & Ratanamahatana, 2005). For two trajectories $x=\{x_{i}\}_{i\in\{1,...,t_{1}\}}$, $\bar{x}=\{\bar{x}_{i}\}_{i\in\{1,...,t_{2}\}}$ and some distance function $d(\cdot,\cdot)$, let DTWD be
$\textbf{DTWD}(x,\bar{x})$ $=\frac{1}{t_{1}}\sum_{i=1}^{t_{1}}\Big{(}\min_{\bar{x}_{j}\in\bar{x}}d(x_{i},\bar{x}_{j})\Big{)}$
$\qquad+\frac{1}{t_{2}}\sum_{i=1}^{t_{2}}\Big{(}\min_{x_{j}\in x}d(\bar{x}_{i},x_{j})\Big{)}$ (21)
4.3 Model Details
@@
As the authors of NCDS have not yet made their code publically available, we implemented NCDS to the best of our abilities. We used a latent dimension of size 2 for all datasets. This is in-line with the description in (Beik-Mohammadi et al., 2024), and necessary given the extra cost from integrating the Jacobian in higher-dimensional latent spaces. We use the same encoder $\phi$ as for ELCD with an added un-padding layer as the last function which removes the last $d-2$ dimensions.
LASA-2D LASA-4D LASA-8D
NCDS 0.59 $\pm$ 0.61 2.19 $\pm$ 1.23 5.07 $\pm$ 1.32
ELCD 0.46 $\pm$ 0.57 1.34 $\pm$ 0.95 2.78 $\pm$ 0.05
Table 1: Average and standard deviation of DTWD between learned trajectories and reference trajectories from 2D, 4D, and 8D-LASA datasets.
4.4 Results
Table 1 presents our results. ELCD performs better than NCDS on average on LASA-2D, 4D, and 8D. With the improved results more pronounced in higher dimensions. We additionally report the mean time per epoch across all data in table (2), which shows ELCD is far more computationally efficient.
LASA-2D LASA-4D LASA-8D
NCDS 73 93 174
ELCD 12 36 108
Table 2: Average time (in seconds) per epoch of training on a single curve (three trajectories) from the 2D, 4D, and 8D-LASA datasets.
5 Conclusions
In this paper, we introduce ELCD, the first neural network-based dynamical system with global contractivity guarantees in arbitrary metrics. The main theoretical tools are extended linearization, equilibrium contraction, and a converse contraction theorem. To allow for contraction with respect to more general metrics, we use a latent space representation with dimension of the latent space equal to the dimension of the data space and train diffeomorphisms between these spaces. We highlight key advantages of ELCD compared to NCDS as introduced in (Beik-Mohammadi et al., 2024), including global contraction guarantees, expressible parameterization, and efficient inference. We demonstrate the performance of ELCD on the 2D, 4D, and 8D-LASA datasets and show significant performance improvement compared to NCDS.
Acknowldegments
This work was supported in part by the NSF Graduate Research Fellowship under grant 213931, the AFOSR under award FA9550- 22-1-0059. This material is additionally based upon work supported by the NSF under grant no. 2229876 and is supported in part by funds provided by the NSF, by the Department of Homeland Security, and by IBM. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the NSF or its federal agency and industry partners.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning with specific applications for dynamical systems and robotics. There are many potential societal consequences of our work, none of which are unique to it among other works in the area. We do not feel any of these consequences need to be specifically highlighted here.

### v3

@@ 3.5 Choice of Diffeomorphism
There are several popular neural network diffeopmorphisms including coupling layers [11, 29], normalizing flows [35], $\mathcal{M}$-flow [8, 5], spline flows [13], and radial basis functions [6].
A coupling layer $\phi:{}^{d}\to{}^{d}$ consists of a neural network $\theta:{}^{k}\to{}^{N}$ with $1<k<d$ and a neural network $g:{}^{N}\times\real\to\real$. The transform $\phi$ maps input $x\in{}^{d}$ to $y\in{}^{d}$ with the following procedure:
@@
Polynomial spline curves are another diffeomorphism that are commonly used as $g$ in coupling layers. A polynomial spline has a restricted domain which is divided into bins. A different linear, quadratic [28], or cubic [12] polynomial covers each bin. [13] introduce rational-quadratic splines, which constructs a spline from functions that are the quotient of two quadratic polynomials.
In practice, $\phi$ is the composition of several of transforms. Because the coupling transform only alters some coordinates, they are often used in conjunction with random permutations. If the latent space dimension is smaller than the data space dimension [5] makes the last composite function of $\phi$ an ’Unpad’ function, which simply removes dimensions to match the latent space. The decoder $\phi$ is then prepended by a ’Pad’ function, which concatenates the latent variable with the necessary number of ’zeros’ to match the data space.
Of course, if the latent dimension is smaller than that of the data dimension, then $\phi$ can not be bijective. However, [5] argue that as long as $\phi$ is injective and has range over the dataset, then $f$ being contracting in the latent space implies that the learned dynamics are contracting on the submanifold defined by the image of $\phi^{-1}$. In other words, NCDS cannot be globally contracting. The consequences of this are shown in Figure (2). While NCDS can learn to admit an equilibrium point when on the submanifold, trajectories that fall off the submanifold may diverge. ELCD, in contrast, is contracting to the correct equilibrium point during all phases of training.
Figure 2: Plots of the vector fields and induced trajectories learned by ELCD (Top) and NCDS (Bottom) after different training epochs. ELCD is contracting always while NCDS may admit multiple equilibrium or diverge.
@@
3.6 Training
Our task is to simultaneously learn the contracting system $f(x)$ and the metric in which $f$ contracts. As previously discussed, the metric is implicitly determined by $\phi(x)$. [5] uses a two-step method. First, they treat $\phi$ as a variational autoencoder and minimize the evidence lower bound (ELBO). They then fix the encoder and train the model to evolve the state in latent space. Note that the VAE objective is to make the encoded data to match a standard-normal distribution. Notably, VAE training does not encourage the encoder to transform the data to space in which the data is corresponds to trajectories of a contracting system. If the data is not contracting in the transformed space, a contracting model will not be able to fully fit the data. This explains why, in our implementation of the two step-training, we are unable to reasonably learn the data. For further comparison with NCDS, we will train the encoder and model jointly.
@@ 4 Experiments
For all datasets we compare our method against NCDS  [5], Stable Deep Dynamics (SDD) [27] and Euclideanizing Flow (Eflow) [29]. See A.2 for a detailed discussion of the methods. We report the dynamic time warping distance (DTWD) [19] (see A.1) and standard deviation between predicted and data trajectories. See A.4 for model implementation details. See also the Appendix for more details on these models.
@@ 4.1 Datasets
We experiment with the LASA dataset [24], which consists of 30, two-dimensional curves. We use three demonstration trajectories of each curve. As in [5], the first few initial points of each trajectory are omitted so that only the target state has zero velocity. we stack two and four LASA trajectories together to make datasets of 4 and 8-dimensional trajectories, respectively. All data is standardized to have a mean of zero and variance of one. We use in total 10 2D curves, 6 4D curves, and 6 8D curves. Some 2D-LASA trajectories and their respective trained ELCD trajectories and vector fields are visualized in Figure (3).
We also experiment with simulated datasets. We simulate 6 trajectories of a 2,4, and 8-link pendulum (4D, 8D, and 16D respectively) each and 4 trajectories of 8D and 16D Riemannian gradient descent dynamics on a generalization of the Rosenbrock function (see Appendix A.3). Each model is trained on all trajectories of the same dimension, and then predictions are made starting from every initial point.
Table 1: DTWD on LASA, Pendulum, and Rosenbrock datasets
{tblr}
hline1-2,5, 8, 10 = 1-5, & SDD EFlow NCDS ELCD
LASA-2D 0.37 $\pm$ 0.32 1.05 $\pm$ 0.25 0.59 $\pm$ 0.61 0.46 $\pm$0.57
LASA-4D 2.49 $\pm$ 2.4 2.24 $\pm$ 0.12 2.19 $\pm$ 1.23 1.34 $\pm$ 0.95
LASA-8D 5.26 $\pm$ 0.50 2.66 $\pm$ 0.63 5.04 $\pm$ 0.77 2.41 $\pm$ 0.68
Pendulum-4D 0.49 $\pm$ 0.11 0.17 $\pm$ 0.01 1.35 $\pm$ 2.260.16 $\pm$ 0.05
Pendulum-8D 0.75 $\pm$ 0.08 0.33 $\pm$ 0.01 2.88 $\pm$ 0.69 0.28 $\pm$ 0.10
Pendulum-16D 1.86 $\pm$ 0.14 0.45 $\pm$ 0.01 1.65 $\pm$ 0.31 0.51 $\pm$ 0.13
Rosenbrock-8D NaN 1.90 $\pm$ 0.16 2.74 $\pm$ 0.15 1.24 $\pm$ 0.26
Rosenbrock-16D NaN 3.57 $\pm$ 0.66 3.68 $\pm$ 0.12 2.26 $\pm$ 0.51
4.2 Results
Table 1 presents our results. ELCD performs the best in all but two datasets, where it performs second-best. These results shows the benefit of the increased expressiveness allowed by the skew symmetric component of our parametrization. EFlow does perform comparable to ELCD on the pendulum datasets, slightly beating us for the 8-link, 16-dimensional setting. As Eflow represents its vector field as the gradient of a function, it is arguably well-suited for the modeling of a mechanical system, which is inherently gradient like due to the underlying energy function.
The Rosenbrock dynamics are stiff and difficult to learn. In our training of SDD, we observed lack of stability and convergence, resulting in very large DTWDs. Hence, we report NaN for the results of the SDD models on the Rosenbrock dataset.
5 Limitations
Our method assumes that the underlying dynamics of the data trajectories are contracting. However, the diffeomorphic transform allows for a wide class of stable systems to be represented. Our method is currently implemented for trajectories that converge to the same fixed point. But, this can be overcome by manually changing the fixed point, depending on which trajectory the input data came from. Also, right now our method is limited to dynamics in n, but we imagine that an extension to more general manifolds should be possible.
6 Conclusions
In this paper, we introduce ELCD, the first neural network-based dynamical system with global contractivity guarantees in arbitrary metrics. The main theoretical tools are extended linearization, equilibrium contraction, and a converse contraction theorem. To allow for contraction with respect to more general metrics, we use a latent space representation with dimension of the latent space equal to the dimension of the data space and train diffeomorphisms between these spaces. We highlight key advantages of ELCD compared to NCDS as introduced in [5], including global contraction guarantees, expressible parameterization, and efficient inference. We demonstrate the performance of ELCD on high-dimensional LASA datasets and simulated Pendulum and Rosenbrock dynamics. Our method shows consistent performance across all datasets.
Acknowledgments and Disclosure of Funding
This work was supported in part by the NSF Graduate Research Fellowship under grant 213931, the AFOSR under award FA9550- 22-1-0059, the NSF under grant no. 2229876, the Department of Homeland Security, and by IBM. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the NSF or its federal agency and industry partners.
Appendix A Appendix
A.1 Metric
@@
We evaluate trajectories by comparing them against their reference trajectory using the normalized dynamic time warping distance (DTWD) [19]. For two trajectories $x=\{x_{i}\}_{i\in\{1,...,t_{1}\}}$, $\bar{x}=\{\bar{x}_{i}\}_{i\in\{1,...,t_{2}\}}$ and some distance function $d(\cdot,\cdot)$, let DTWD be
$\textbf{DTWD}(x,\bar{x})$ $=\frac{1}{t_{1}}\sum_{i=1}^{t_{1}}\Big{(}\min_{\bar{x}_{j}\in\bar{x}}d(x_{i},\bar{x}_{j})\Big{)}+\frac{1}{t_{2}}\sum_{i=1}^{t_{2}}\Big{(}\min_{x_{j}\in x}d(\bar{x}_{i},x_{j})\Big{)}$
A.2 Comparisons to Other Models
In this section, we describe the models that we compare to, namely Eflow [29], SDD [27], and NCDS [5].
Euclideanizing Flow: Starting with Eflow, they parametrize their dynamical system by
$\dot{x}=-G_{\psi}(x)^{-1}\nabla\Phi(\psi(x)),$ (18)
where $\psi:{}^{d}\rightarrow{}^{d}$ is a diffeomorphism, $\Phi:{}^{d}\rightarrow\real$ is a convex potential function, and $G_{\psi}(x)=D\mkern-0.75mu{\psi}(x)^{\top}D\mkern-0.75mu{\psi}(x)$ is the induced Riemmanian metrc. The dynamics (18) then has the interpretation of being the steepest descent for $\Phi\circ\psi$ on the Riemmanian manifold defined by metric $G_{\psi}$. Specifically, in [29], the convex potential function is defined to be $\Phi(y)=\|y-y^{\star}\|_{2}$, where $y^{\star}=\psi(x^{\star})$ and the diffeomorphism is parametrized via a coupling layer [11]. Note that contraction properties of dynamics of these form were studied in [37]. Since $\Psi$ is chosen to be convex but not strongly convex, these dynamics are only weakly contracting, i.e., satisfy (2) with $c=0$.
Stable Deep Dynamics: In [27], the authors parametrize both an unconstrained vector field $\hat{f}:{}^{d}\rightarrow{}^{d}$ and a Lyapunov function $V:{}^{d}\rightarrow\mathbb{R}_{\geq 0}$ via neural networks. Then to enforce global exponential stability, at every point $x\in{}^{d}$, they project $\hat{f}(x)$ onto the convex set of points $\{u\in{}^{d}\;|\;\nabla V(x)^{\top}u\leq-\alpha V(x)\}$. They do this projection in closed-form so that the dynamics have the representation
$\dot{x}=\hat{f}(x)-\nabla V(x)\frac{\mathrm{ReLU}(\nabla V(x)^{\top}\hat{f}(x)+\alpha V(x))}{\|\nabla V(x)\|_{2}^{2}}.$ (19)
Under a smart parametrization of $V$ using input-convex neural networks [3], the authors guarantee global exponential stability of (19) to the origin with rate $\alpha$. Note that the dynamics (19) are continuous, but not necessarily differentiable. Due to this potential nonsmoothness, it is theoretically unknown whether these dynamics are contracting (since Theorem 2.8 in [15] requires at least some differentiability).
Neural Contractive Dynamical Systems: In [5], the authors enforce contractivity by directly enforcing negative definiteness of the Jacobian matrix of the vector field by parametrizing
$D\mkern-0.75mu{f}(x)=-(J_{\theta}(x)^{\top}J_{\theta}(x)+\epsilon I_{d}),$ (20)
where $J_{\theta}:{}^{d}\rightarrow{}^{d\times d}$ is a neural network. Under this parametrization of the Jacobian, it is straightforward to see that (2) holds with $c=\epsilon$ and $M(x)=I_{d}$ for all $x$. To recover the vector field from the Jacobian, the fundamental theorem of calculus for line integrals (which can be seen as a version of the mean-value theorem) is utilized to express
[…]
where $x_{0}$ and $f(x_{0})$ denote the initial condition and its velocity, respectively.
To express richer dynamics, NCDS also consists of an encoder and decoder for a lower-dimensional latent space. The dynamics in the latent space are constrained to be (21) and the encoder and decoder are parametrized using a VAE. Since, generally, the latent space is of lower dimension than the data space, the resulting dynamics are only guaranteed to be contracting on the submanifold defined by the image of the decoder.
A.3 Additional Details on Datasets
In this section, we provide additional elaboration on the $n$-link pendulum dataset and the Rosenbrock dataset.
$n$-link pendulum: The $n$-link pendulum is a simple mechanical system which is the interconnection of $n$ rigid links under the force of gravity. We specifically assume that there is damping on each of these links, which makes the resulting dynamical system stable and almost all trajectories converge to the downright equilibrium point. In this case, the state of the pendulum is described by the $n$ pairs $(\theta,\dot{\theta}_{i})$, where $\theta$ is the angle of the $i$-th link and $\dot{\theta}_{i}$ is its angular velocity. Thus, this dynamical system is $2n$ dimensional.
As was done in [27], we adapt the code from http://jakevdp.github.io/blog/2017/03/08/triple-pendulum-chaos/ to generate data for $n$ links.
Rosenbrock datset: The classical Rosenbrock function, see [31]
$f(x,y)=(1-x)^{2}+100(y-x^{2})^{2},$ (22)
is a nonconvex function with global minimum at $(1,1)$. Despite its apparent nonconvexity, it is known that a Riemannian gradient descent dynamics, is contracting with respect to a suitable Riemannian metric, see Example 3 in [37] for details. In optimization and in machine learning, the Rosenbrock function is a benchmark for the design of various optimizers and in the study of Riemannian optimization [18, 23].
For a higher-dimensional generalization of the Rosenbrock function, we consider
$f(x)=\lambda_{1}(1-x_{1})^{2}+\sum_{i=2}^{n}\lambda_{i}(x_{i}-x_{i-1}^{2})^{2},$ (23)
where $\lambda_{i}>0$ are parameters that affect the conditioning of the function. This generalization is still nonconvex and admits several saddle points. Note that the classic Rosenbrock corresponds to $n=2,\lambda_{1}=1,\lambda_{2}=100$. The unique global minimum of this generalization is at $(1,1,\dots,1)$. The corresponding contracting Riemmanian gradient descent dynamics that find this global minimum are
$\dot{x}=-G(x)^{-1}\nabla f(x),$ (24)
where $G(x)=D\mkern-0.75mu{\psi}(x)^{\top}D\mkern-0.75mu{\psi}(x)$, where $\psi:{}^{n}\rightarrow{}^{n}$ is the mapping $\psi(x)=(\sqrt{\lambda_{1}}(1-x_{1}),\sqrt{\lambda_{2}}(x_{2}-x_{1}^{2}),\dots,\sqrt{\lambda_{n}}(x_{n}-x_{n-1}^{2}))$. Note that these dynamics were designed via a generalization of the procedure proposed in [37]. To generate data, we choose four initial points and evolved them according to  (24).
A.4 Model Details
@@
As the authors of NCDS have not yet made their code publicly available, we implemented NCDS to the best of our abilities. We used a latent dimension of size 2 for all datasets. This is in-line with the description in [5], and necessary given the extra cost from integrating the Jacobian in higher-dimensional latent spaces. We use the same encoder $\phi$ as for ELCD with an added un-padding layer as the last function which removes the last $d-2$ dimensions.
A.5 Training details
All experiments are trained with a batch size of 100, for 100 epochs, with an Adam optimizer and learning rate of 1e-3. All computation is done on CPUs.